A buffer solution contains 0.59 mol of ascorbic acid (HC6H7O6) and 0.30 mol of sodium ascorbate (NaC6H7O6) in 3.70 L. The Ka of ascorbic acid (HC6H7O6) is Ka = 8e-05.

(a) What is the pH of this buffer? pH =

(b) What is the pH of the buffer after the addition of 0.28 mol of NaOH? (assume no volume change) pH =

(c) What is the pH of the original buffer after the addition of 0.11 mol of HI? (assume no volume change) pH =
Answer #1

A) it is acidic buffer

pKa = 4.10

pH = pKa + log(Salt/acid)

=4.10+log(0.30/0.59)

pH = 3.80

B) if we add NaOH it reacts with Acid

So, acid remains = 0.59-0.28 = 0.31 mol

Mole of salt formed = 0.30+.28 =0.58 mol

pH = 4.10 + log(0.58/0.31)

=> 4.37

C) if we add HI, it reacts with conjugate base

So acid formed = 0.11 + 0. 59 = 0.70 mol

Conjugate base remains= 0.30-0.11=> 0.19 mol

pH = 4.10+(0.19/0.70)

=> 3.53